OpenStudy (anonymous):

Nadia rented a stage platform at $410 for 4 days. If she rents the same stage platform for 8 days, she has to pay a total rent of $770. Write an equation in the standard form to represent the total rent (y) that Nadia has to pay for renting the stage platform for x days Please HELP!!.

OpenStudy (anonymous):

are you learning about exponents in your class? does y=ab^x sound familiar?

OpenStudy (anonymous):

@MathLegend they leave when I try to help, lol

OpenStudy (mathlegend):

I never learned y=ab^x can you further explain this @DemolisionWolf ?

OpenStudy (anonymous):

No I haven't heard of that...

OpenStudy (anonymous):

@mathlegend yep, so y=ab^x is the standard form to write a formula that reflects an exponential function. so, looking at a table made from the two 'points' given we have: D | $ 4 410 8 770 so we can tell that the relationship from one day to another day is not the same, so the function isn't linear, so i assumed it's going to be an exponent (but now looking, kaileam posted that it's not an exponential function, but for the sake of teaching you, i'll keep going) so if we let x= Day and y=$cost then we can get a system of equations from the two known points. 410 = ab^4 770 = ab^8 we can solve these two equation with substitution and get 770 = (410/b^4)b^8 simplify 770/410 = b^4 b=1.17 back substitue 770=a(1.17^8) a=219.28 so, had this question been about exponential functions, the solution function would have been y=(219.28)1.17^x make sense?

OpenStudy (anonymous):

@kaileam are you guys learning about 'linear functions' or y=mx+b?

OpenStudy (anonymous):

Yes I have learned that just not what you mentioned above

OpenStudy (anonymous):

ok ^_^

OpenStudy (anonymous):

@MathLegend I jumped to conclusions too quickly when i first asked if it was an exponential function, because i'm so excited to use that formual I just learned last week, lol

OpenStudy (anonymous):

I get that the question is asking to be written is standard form: Ax + By = C, it;s just that I don't know how to do that in this type of situation

OpenStudy (anonymous):

@kaileam i'll show you ^_^ we'll want to use the y=mx+b to figure this question out. so lets say that days = x and cost = y then we would have the points (4,410) and (8,770) with these two points, we can first solve for the slope, then solve for b

OpenStudy (anonymous):

@MathLegend I made a slip up, i said in my post to you, 'we can see that the function is not linear' was wrong.

OpenStudy (anonymous):

okay...

OpenStudy (anonymous):

The slope is 90

OpenStudy (anonymous):

can you find the slope between the two points?

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

sorry! my computer was waaaaaay slow

OpenStudy (anonymous):

so it would be y = 90x + 50

OpenStudy (anonymous):

same here! I think it was a glitch

OpenStudy (anonymous):

so then the final equation in standard form would be 90x - y = -50 right?

OpenStudy (anonymous):

yep ^_^ good work

OpenStudy (anonymous):

Thanks a lot! I really appreciate it!

OpenStudy (anonymous):

ur welcome hvae a good day ^_^

OpenStudy (anonymous):

you too

OpenStudy (mathlegend):

@DemolisionWolf shouldn't he have gotten -90x + y= 50

OpenStudy (anonymous):

@MathLegend ya... they are the same equation, i'm not to knowledgeable on proper notation tho...

OpenStudy (mathlegend):

Yeah but @kaileam has it incorrectly

OpenStudy (mathlegend):

incorrect

OpenStudy (mathlegend):

Wrong signs